# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v3

@@ A.4 Lemmas for Utility Analysis
Lemma A.14 (Implications of smoothness (beck2017first)).
@@
Lemma A.15 (Implications of strong convexity (beck2017first)).
@@ Appendix C Proofs for Utility Analysis
Our utility analysis builds upon the results of bassily2014private, zhang2020improved, and koloskova2023revisiting.
@@ C.2 Privacy-utility Trade-off for DPSGD-GC
Implications. Prior work (see, e.g., (bassily2014private), (wang2017differentially), and (zhang2017efficient)) have demonstrated that the optimal utility of $(\epsilon,\delta)$-DP algorithms on Lipschitz smooth population risk functions with bounded stochastic gradient assumption is $\frac{\sqrt{dL\log(1/\delta)}}{\epsilon n}$. In the case of small stochastic noise, we achieve a utility bound of $\frac{dL\log(1/\delta)}{\epsilon^{2}n^{2}}$, which is of higher order—i.e., slightly worse than the optimal utility—due to the relaxed assumption of bounded gradients and a different utility metric. When the stochastic noise scale $\sigma_{\text{SGD}}$ is large, the error introduced by the SGD variance becomes dominant and can not be effectively reduced-consistent with our intuition.
@@ C.3 Privacy-utility Trade-off for DPSGD-DC
Prior work (see, e.g., (bassily2014private), (wang2017differentially), and (zhang2017efficient)) have also demonstrated that the optimal utility of $(\epsilon,\delta)$-DP algorithms on Lipschitz smooth and strongly convex population risk functions with bounded stochastic gradient assumption is $\frac{dL\log(1/\delta)}{\mu\epsilon^{2}n^{2}}$. In the case of small stochastic noise and small total iterations $T$, our resulting utility bound matches the theoretical bound—up to a constant—for the best attainable utility. For other cases, the error introduced by the SGD variance and domain norm becomes dominant and can not be effectively reduced-consistent with our intuition.
@@ D.2 MIA Setting
We train a private ResNet-18 network (he2016deep) as the target model using the Opacus library (yousefpour2021opacus) on the standard CIFAR10 dataset (krizhevsky2009learning). According to standard MIA protocols (shokri2017membership), the training and test sets for each target and shadow model are randomly selected, equal in size, and mutually disjoint. We set the training set size to $10,000$. While the target model’s dataset does not overlap with those of the shadow models, different shadow models may partially share the same data. The shadow models are trained using the same architectures as the target model. For the attack model, we employ a two-layer multilayer perceptron (MLP) with $50$ hidden nodes and ReLU as activation functions.
@@ D.4 Implementation Details
For the target model, we use the SGD optimizer with the learning rate $\eta=0.1$, the noise level $\sigma_{\text{DP}}=0.002$, the clipping threshold $C=20$, and the confidence level $\delta=10^{-5}$. We train $10$ shadow models to simulate the behavior of the target model. For the attack classifier, we use the SGD optimizer with the initial learning rate of $0.01$, the weight decay of $5\times 10^{-4}$, and the momentum of $0.9$. The mini-batch size during training is set to be $100$. We predict the labels by selecting from the model’s output the class with the highest probability. Experimental results are reported by averaging over $10$ Monte Carlo trials. Our experiments are implemented in the PyTorch (paszke2017automatic) framework.
@@ E.1 Discussion of Bounded Domain Assumption
We provide a detailed discussion of the bounded domain assumption below. Note that a naive yet convergent bound can be obtained by combining the post-processing property (Lemma A.9) and the standard Gaussian mechanism sensitivity analysis (dwork2014algorithmic) under a bounded domain assumption. For instance, one such naive bound can be easily derived as: $\frac{2\alpha(D+\frac{\eta}{b}C)^{2}}{\sigma_{\text{DP}}^{2}}$. However, this simplistic bound severely overestimates the privacy loss, leading to overly conservative noise requirements and consequently deteriorating the privacy-utility trade-off. In this paper, we rigorously derive a tighter, non-trivial convergent privacy bound. To highlight this, we have compared our derived bound against this trivial bound in Figure 5. For additional context, recent works such as (altschuler2022privacy), (kong2024privacy), and (chien2024convergent) have also adopted the bounded domain assumption to achieve convergent guarantees.
Also, we emphasize that the bounded domain assumption is mild and practically justified. Many optimization problems naturally operate within constrained parameter domains due to intrinsic problem characteristics or practical requirements. Furthermore, some unconstrained problems may be effectively addressed by solving a sequence of constrained sub-problems, which incurs only a minor computation and privacy overhead (liu2019private). Consequently, this assumption is widely recognized as mild, practical, and standard in both optimization theory and privacy-preserving algorithmic design.
@@ E.2 Discussion of Smoothness Assumption
The smoothness parameter, $L$, can be estimated numerically, e.g., as the largest eigenvalue of the data Gram matrix for convex losses. While our utility analysis can be extended to broader smoothness (for example, $(L_{0},L_{1})$-smooth (koloskova2023revisiting)), our RDP analysis requires the smooth assumption, and may generalize to Hölder continuous gradients (chien2024convergent).

### v4

@@ A.4 Lemmas for Utility Analysis
Lemma A.14 (Implications of smoothness (Beck 2017)).
@@
Lemma A.15 (Implications of strong convexity (Beck 2017)).
@@ Appendix C Proofs for Utility Analysis
Our utility analysis builds upon the results of Bassily et al. (2014), Zhang et al. (2020), and Koloskova et al. (2023).
@@ C.2 Privacy-utility Trade-off for DPSGD-GC
Implications. Prior work (see, e.g., (Bassily et al. 2014), (Wang et al. 2017), and (Zhang et al. 2017)) have demonstrated that the optimal utility of $(\epsilon,\delta)$-DP algorithms on Lipschitz smooth population risk functions with bounded stochastic gradient assumption is $\frac{\sqrt{dL\log(1/\delta)}}{\epsilon n}$. In the case of small stochastic noise, we achieve a utility bound of $\frac{dL\log(1/\delta)}{\epsilon^{2}n^{2}}$, which is of higher order—i.e., slightly worse than the optimal utility—due to the relaxed assumption of bounded gradients and a different utility metric. When the stochastic noise scale $\sigma_{\text{SGD}}$ is large, the error introduced by the SGD variance becomes dominant and can not be effectively reduced-consistent with our intuition.
@@ C.3 Privacy-utility Trade-off for DPSGD-DC
Prior work (see, e.g., (Bassily et al. 2014), (Wang et al. 2017), and (Zhang et al. 2017)) have also demonstrated that the optimal utility of $(\epsilon,\delta)$-DP algorithms on Lipschitz smooth and strongly convex population risk functions with bounded stochastic gradient assumption is $\frac{dL\log(1/\delta)}{\mu\epsilon^{2}n^{2}}$. In the case of small stochastic noise and small total iterations $T$, our resulting utility bound matches the theoretical bound—up to a constant—for the best attainable utility. For other cases, the error introduced by the SGD variance and domain norm becomes dominant and can not be effectively reduced-consistent with our intuition.
@@ D.2 MIA Setting
We train a private ResNet-18 network (He et al. 2016) as the target model using the Opacus library (Yousefpour et al. 2021) on the standard CIFAR10 dataset (Krizhevsky and Hinton 2009). According to standard MIA protocols (Shokri et al. 2017), the training and test sets for each target and shadow model are randomly selected, equal in size, and mutually disjoint. We set the training set size to $10,000$. While the target model’s dataset does not overlap with those of the shadow models, different shadow models may partially share the same data. The shadow models are trained using the same architectures as the target model. For the attack model, we employ a two-layer multilayer perceptron (MLP) with $50$ hidden nodes and ReLU as activation functions.
@@ D.4 Implementation Details
For the target model, we use the SGD optimizer with the learning rate $\eta=0.1$, the noise level $\sigma_{\text{DP}}=0.002$, the clipping threshold $C=20$, and the confidence level $\delta=10^{-5}$. We train $10$ shadow models to simulate the behavior of the target model. For the attack classifier, we use the SGD optimizer with the initial learning rate of $0.01$, the weight decay of $5\times 10^{-4}$, and the momentum of $0.9$. The mini-batch size during training is set to be $100$. We predict the labels by selecting from the model’s output the class with the highest probability. Experimental results are reported by averaging over $10$ Monte Carlo trials. Our experiments are implemented in the PyTorch (Paszke et al. 2017) framework.
@@ E.1 Discussion of Bounded Domain Assumption
We provide a detailed discussion of the bounded domain assumption below. Note that a naive yet convergent bound can be obtained by combining the post-processing property (Lemma A.9) and the standard Gaussian mechanism sensitivity analysis (Dwork et al. 2014) under a bounded domain assumption. For instance, one such naive bound can be easily derived as: $\frac{2\alpha(D+\frac{\eta}{b}C)^{2}}{\sigma_{\text{DP}}^{2}}$. However, this simplistic bound severely overestimates the privacy loss, leading to overly conservative noise requirements and consequently deteriorating the privacy-utility trade-off. In this paper, we rigorously derive a tighter, non-trivial convergent privacy bound. To highlight this, we have compared our derived bound against this trivial bound in Figure 5. For additional context, recent works such as (Altschuler and Talwar 2022), (Kong and Ribero 2024), and (Chien and Li 2024) have also adopted the bounded domain assumption to achieve convergent guarantees.
Also, we emphasize that the bounded domain assumption is mild and practically justified. Many optimization problems naturally operate within constrained parameter domains due to intrinsic problem characteristics or practical requirements. Furthermore, some unconstrained problems may be effectively addressed by solving a sequence of constrained sub-problems, which incurs only a minor computation and privacy overhead (Liu and Talwar 2019). Consequently, this assumption is widely recognized as mild, practical, and standard in both optimization theory and privacy-preserving algorithmic design.
@@ E.2 Discussion of Smoothness Assumption
The smoothness parameter, $L$, can be estimated numerically, e.g., as the largest eigenvalue of the data Gram matrix for convex losses. While our utility analysis can be extended to broader smoothness (for example, $(L_{0},L_{1})$-smooth (Koloskova et al. 2023)), our RDP analysis requires the smooth assumption, and may generalize to Hölder continuous gradients (Chien and Li 2024).